Susan Marciano invested part of her 16,000 bonus in a fund that paid an 8% profit and invested the rest in stock that suffered a 5% loss. Find the amount of each investment if her overall net profit was 890.
Answer by addingup(3677)

Amount invested at 8%: x
Amount invested at -5%: 16,000-x
0.08x-0.05(16,000-x) = 890
0.08x-800+0.05x = 890
0.13x = 1690
x = 13,000
<><><><><><><><><><><><><><>
Check:
at 8%: 13,000*0.08 = 1,040
at 5% (16,000-13,000)*0.05 = 150
Subtract, since the amount at 5% was a loss:
1,040-150 = 890 Correct
:
John